OpenStudy (anonymous):

if 2/3 p + 6 =7/6 p , what is p?

OpenStudy (lynfran):

bring p to 1side of the equation and solve

OpenStudy (lynfran):

hint: subtract 2/3p from both side of the equation

OpenStudy (lynfran):

so we have 6=7/6p-2/3p

OpenStudy (jdoe0001):

hint: multiply both sides by "p"

OpenStudy (anonymous):

r u guys sure?

OpenStudy (anonymous):

@jdoe0001 u r wrong

OpenStudy (anonymous):

if we follow what @lynfran stated we can arrive at the answer

OpenStudy (jdoe0001):

hmm? is the fellow even here? he doesn't seem to be saying anything

OpenStudy (anonymous):

idk

OpenStudy (lynfran):

\[\frac{ 2 }{ 3p }+6=\frac{ 7 }{ 6p }\]\[6=\frac{ 7 }{ 6p }-\frac{ 2 }{ 3p }\]\[6=\frac{ 7-4 }{ 6p }\]\[6=\frac{ 3 }{ 6p }\]\[6=\frac{ 1 }{ 2p }\]multiply both side by 2p we get \[2p*6=1\]\...solve for p..

OpenStudy (anonymous):

@LynFran woudnt it be like this? \[\frac{ 2 }{ 3 }p+6=\frac{ 7 }{ 6 }\] \[6=\frac{ 7 }{ 6 }p-\frac{ 2 }{ 3 }p\] (you subtract 7/6 - 2/3 and get 1/2 ) \[6=\frac{ 1 }{ 2 }p\] \[p=?\]

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

no... then u will have p = 12 thats not the answer...lol

OpenStudy (anonymous):

no?

OpenStudy (lynfran):

continue frm where i left off 2p*6=1 12p=1 divide both sides by 12 to get the value of p...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but that would not check out right. you would get, if you plug in 1/12 in for p, 109/18 = 7/72. wheras if you plug in 12 (the answer that i got ) you would get 14 = 14 clearly checking out.

OpenStudy (anonymous):

or am i missing something?

OpenStudy (lynfran):

\[\frac{ 2 }{ 3(\frac{ 1 }{ 12 }) }+6=?\]

OpenStudy (anonymous):

14

OpenStudy (anonymous):

what ???

OpenStudy (lynfran):

\[\frac{ 7 }{ 6(\frac{ 1 }{ 12 }) }=?\]

OpenStudy (anonymous):

ok idk how you did that

OpenStudy (lynfran):

lol..

OpenStudy (anonymous):

but then woudnt that mean that im also right?

OpenStudy (anonymous):

couse idk what i did wrong

OpenStudy (lynfran):

ok lets check...\[\frac{ 2 }{ 3(12) }+6=?\]\[\frac{ 7 }{ 6(12) }=?\]

OpenStudy (lynfran):

u was suppose to find a common denominator \[\frac{ 7 }{ 6p }-\frac{ 2 }{ 3p }\]

OpenStudy (anonymous):

wasnt the question \[\frac{ 2 }{ 3 }P+6=\frac{ 7 }{ 6 }P\] but you solved for \[\frac{ 2 }{ 3P }+6=\frac{ 7 }{ 6p }\]

OpenStudy (anonymous):

isnt there a differance?

OpenStudy (lynfran):

so \[\frac{ (7*1)-(2*2) }{ 6p }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2p }\]

OpenStudy (lynfran):

and remember p" is in the denominator so we multiply by 2p to both sides of equation

OpenStudy (lynfran):

so 2p*6=1

OpenStudy (anonymous):

but in the original problem how do we know that the p was on the denominator?

OpenStudy (lynfran):

well it looked so to me lol...could be wrong then...

OpenStudy (lynfran):

if its not then ur right..

OpenStudy (anonymous):

i understand what you were doing tho

OpenStudy (lynfran):

ok..

OpenStudy (lynfran):

but tylerhuang is offline hmm...